User Tools

Site Tools


List of problems found

Active general issues


2019 Session 3

L - Band

Jodrell Bank (Lovell):

There is no ANTABFS files produced from station for any of the Lovell L band observations, due to a new source of RFI identified at Jodrell Bank which it appears has had very deleterious effects on all the Lovell L band observations for this session. This new RFI is approximately 30 dB above the noise and 20 Mega Hertz wide, spanning 1472 - 1492 Mega Hertz.

The RFI is believed to be for a 4G+ mobile phone service which will bridge the gap between 4G and 5G. It is thought that the signal provides coverage to the Manchester metropolitan area, and this is seen at Pickmere and Darnall too.

This signal was first observed in September / October 2019 and they are still devising methods to mitigate its effects.

2019 Session 2

L - Band

Tianma: Recorded linear polarization (eg.user experiments - EG100C-D and EH036A-F).

2019 Session 1

Urumqi: No ANTAB files produced (invalid TPI values recorded for all experiments).
Tianma: In this session they have fixed the high leakage observed in the last year.
Jodrell Bank: Wiggles in LCP across the band since ~2017 session 2 from unknown origin. It is visible in NMEs or spectral line experiments, although in continuum ones the wiggles are smeared out and only a larger scatter is visible.

Last checks from Eskil point to: “One of the receiver boards in one optical-to-RF converter used for Mk2 VLBI seems to be suboptimal. With very high input power, it works, but when decreasing the power to usual or weaker than normal observing level, the jagged bandpass appears. We'll investigate buying a replacement model. In the mean time, we'll use the identical Lovell hardware which appears to suffer less from this problem.” Solved for 2019 Session 2.

L - Band

Sardinia: LL signal in Subband 8 much lower than RR signal.
Svetloe: LCP is broken (at least in the NME).
Urumqi: recorded linear polarization (in the NME and also in the user experiments, eg. EP106E, EP106F and EP106G).
Tianma: recorded linear polarization (in the NME and also in the user experiments, eg. EP106F and EP106G).

C - band

Svetloe: LCP is broken (at least in the NME).

K - band

KVN: high invalid data percentage (~7%).
Sardinia: No LL signal in subband 7 (and bad sampler stats there too).

M - band

Torun: large cross hand amplitude noticed (~30%).

2018 Session 3


Tianma high leakage between cross and direct polarizations. They should fix that.
They also had recording problems: again they overwrite part of the diskpacks.
e-MERLIN: recording issues. But good fringes and data obtained during several experiments.

L - Band

Sardinia LL signal in Subband 8 much lower than RR signal.
Urumqi Observed with linear polarizations.
Westerbork: high leakage between polarizations.

C - Band

Noto: Out in many experiments due to receiver problems.
Kunming: Fringes only appear in prat of the observations, not full time range. Antenna problems?

2018 Session 2


C - Band

L - Band

Westerbork issues finding fringes in NME → found 7s offset in fila10g. fila10g time seems to be jumping out of sync very quickly (1s/half hour).
Sardinia LL signal in Subband 8 much lower than RR signal.
Zelenchukskaya Tsys values high for Subbands R1 and L1 caused by RFI.

S/X - Band

Westerbork issues with fila10g during NME. Would only send 1 packet per scan. Fixed for first user exp.
Effelsberg issues with fila10g during NME. Would only send 1 packet per scan. Fixed by first user exp.
Noto fringes very weak.

K band

Noto fringes very weak.

2018 Session 1


All EVN Major issue with E-series firmware → amplitude calibration very difficult as slope/jumps across band.

C band

Onsala Broken chip in antenna control unit.

L band

Westerbork L-band receiver problems (no fringes found, unclear reasons).
Tianma recorded a single linear polarization. Fixed for the e-EVN in 7 Feb 2018.

2018 e-EVNs

Westerbork (2018/04/06) the maser broke.
Tianma (2018/02/07, L-band) recorded linear polarizations instead of circulars (continuation of the problem observed in 2017 Session 2). Fixed with Martí-Vidal's program.
Noto (2018/04/04, K-band) had a communication problem between the computers in the control room and was impossible to perform the experiment.

2017 Session 3


Large amplitude issues across band with E-series firmware.
Zc tsys bad. Almost always 999s.
Hartebeesthoek encoder issues (continued).

L band:

Jodrell Bank had a warm receiver.
Tianma sent one linear polarisation only initially, then linears.
Urumqui sent linears.
Torun no fringes ea058b. Unknown reason.
Irbene low power LCP. Jodrell Bank warm receiver.
Westebork had some issues with setting incorrect mark5 mask.
Svetloe missed a few scans due to antenna control issues.

S/X band:

Onsala6 had a warm receiver.
Urumqui reported receiver fault. Also formatter failure.
Kunming issues with maser.

C band:

Urumqi has an odd pattern in its TSYS, first four ~20, last four ~100-150.
Onsala Broken chip in antenna control unit.
Jodrell Bank MkII pointing issues. Appear time/elevation dependant.
Effelsberg FS computer froze.
Yebes reported FS problem. Medicina and Westerbork both reported Fila10g loss of sync.

2017 Session 2


Torun record pointer jumping back to 0 → overwriting experiments.

C band:

Noto first four subbands, widely different clock LCP (weak phases).
Westerbork bad weights ~ 0.6 in EG098B EJ019D (these are back to back)
Zelenchukskaya Consistently has less power in one of the bands (I think RCP). Some phasejumps seen in EG088B also…

L band:

Urumqi bad subband 2 (0 counting) LCP
Tianma - poor polarisation performance. They recorded only one linear polarization and transmitted as RR/LL.
Tianma receiver malfunctioned needed to be repaired and warmed up. Missed some experiments.
Torun recorder skipped back to 0 → overwrote scans 1-208 in EG098A

2017 Session 1


Noto had a warm receiver in the Q-band NME.
KVN Ulsan has consistently had issues transferring data for FTP tests.

Yebes had a broken maser, didn't affect observations directly, the resulting change in DBBC caused a change in fringe amplitude accross the band.

the above is unedited what Jay wrote -- the effect in fringe amp may
be because of vector averaging in time with a higher phase(t) slope --
in the March and April e-VLBI runs, we've noticed that the sign convention
of the daily GPS may have changed again compared to what we had been doing:
clock searches there suggested a rate adjustement that effectively changed
the sign of the rate.  In the e-VLBI, this was a net rate change of ~8 ps/s,
or about a radian in about 4s, which could lead to de-correlation in a
vector-averaged amp(freq) or amp(lag) plot

2016 Session 3


The whole session after the initial two 5cm observations ran with DBBC firmware v105E.
Irbene showed what appears to be large packet loss to the recorder resulting in poor data being recorded. The first disk-pack proved unusable.
Sub-band 5 (BBC 5,13 LSB) had a dip in the center of the band, seen in auto- and cross-correlations.
Urumqi had antenna gear-box maintenance, only observed K band.
No fringes to e-MERLIN stations in NME ftp fringe-tests. JBO still working on obtaining local fringes via their SFXC before sending us the full out-station data (have correlated 3 of 4 user experiments that had out-stations – they seem to have most hope for the last one EM122B, which we're holding to last).
Kunming (thus not yet “EVN”) had fringes in 6cm, 5cm, 3.6cm NMEs.May have had a maser problem in 5cm NME (much increased phase noise).

M band:

Effelsberg had a broken amplifier in LCP for the NME resulting in significantly lower power in LCP channels than in RCP ones. Fixed by the immediately following C-band session.

C band:

Hartebeesthoek had a warm receiver in ER045A (cryo failure); also HA encoder jumping intermittently (but not a large problem).
Jodrell had no fringes in EM121B and EP099A, which were 1Gbps observations immediately following the first 2Gbps observation. Another 2Gbps observation followed these two, then another 1Gbps one, and then a series of 2 or 1Gbps ones – all of these subsequent ones were good.
Noto showed its typical falling-off in amplitude in the upper LCP channels(BBCs 13,15 in DBBC firmware v105E, but this has also occurred the “normal” firmware in which it was BBCs 7,8).
Torun had a issue with a pack in which the record-pointer went back to zero after some experiments had already been recorded – cost 4.5 experiments at C- and L-band.
Yebes receiver fault in ER045A (fixed during experiment).
Zelenchukskaya had consistently low power in RCP channels throughout, and a warm receiver in EM126A. Robledo used a new version of their DVP back-end that enables recording fewer than 32 channels, but no fringes in any observation (correlating a single Ro channel against all channels elsewhere showed some response in the wrong place, but with a null amplitude and a 90-deg phase jump in the middle of the band, suggesting that there was an extra transform going from complex to dual-sideband real “BBCs”).

L band:

Irbene had an NTP error in EM123B that prevented recording.
Jodrell Bank had a problem seemingly limited to L-band observations after the NME in which the fringe walked away in jumps of 512 lags every 10-14s. One difference between the NMEs and user experiments was that the former had 8MHz subbands and the latter 16MHz ones (where 512 lags = 32us).

K band:

Noto cryo failure during NME.
KVN Yonsei did not participate in this session due to re-alignment of antenna panels.
KVN Tamna had a hexapod failure in EP099B.

X band:

Noto had mark5 problems in EP099C

2016 Session 2


Noto several times has RCP at half the power of LCP.
Torun BBC5/USB had a different delay (phase vs. freq) for all experiments this session.

C band:

Westerbork broken hybrid - no good circular pols.
Several telescopes reporting FS crashes.

K band:

Urumqi receiver failure.
Effelsberg swapped pols.
Several telescopes (Hh, Ef, Sr, O8) reported issues with v105E firmware giving bad tsys and bit statistics for several BBCs. Fixed by DBBC re-program.

evn/supportscientists/problems.txt · Last modified: 2020/02/05 10:13 by nair